ln(x − 5) + ln(x + 6) = 1

Answer by ankor@dixie-net.com(22740)

ln(x − 5) + ln(x + 6) = 1
we can write it as a single ln
ln((x-5)(x+6)) = 1
FOIL
ln(x^2+x-30) = 1
antilog,
x^2 + x - 30 = e
x^2 + x - 30 ~ 2.7183
x^2 + x - 30 - 2.7183 = 0
x^2 + x - 32.7183 = 0
using the quadratic formula, the positive solution
x = 5.2418
:
:
Check on calc ln(5.2418-5) + ln(5.2418+6) = .9999946 ~ 1